### v2

# A remark on the number of automorphisms of some algebraic structures

Abstract.
In these notes we look at the following question. Given a category $\mathcal{C}$ of algebraic structure (e.g. the category of groups, monoids, partial groups, …) and a rational $r\in\mathbb{Q}$, does there exists an element $x\in\mathcal{C}$ such that the size of its automorphism group $\text{Aut}_{\mathcal{C}}(x)$ divided by the size of $x$ (whatever that would means) is equal to $r$ To our knowledge, this question was introduced by Tărnăuceanu in the category of groups. Here, we answer positively to this question in the categories of graphs, monoids, partial groups and posets.
Key words and phrases:
automorphisms, graphs, monoids, partial groups, posets
1991 Mathematics Subject Classification:
Primary: 08A35 ; Secondary: 05C25, 20F29
1. Introduction
In a recent paper [8], Tărnăuceanu proved the following nice density result.
Theorem 1.1 ([8], Main Theorem).
The set
$\left\{\frac{|\text{Aut}_{{{\mathcal{G}}roups}}(G)|}{|G|}\;\middle|\;G\text{ afinite group}\right\}$
is dense in $\mathbb{Q}\cap]0,+\infty[$.
His result is actually a bit stronger since he proved that the subset considering only finite abelian groups is dense. From this theorem, he address the following question.
Question.
Does there exists, for every $r\in\mathbb{Q}\cap]0;+\infty[$, a group $G$ such that $\dfrac{|\text{Aut}_{{{\mathcal{G}}roups}}(G)|}{|G|}=r$
We are not trying to answer this question here but we look at the question in two more general contexts: the category ${{\mathcal{M}}onoids}$ of monoids and the category ${{\mathcal{P}}art}$ of partial groups. Theses two algebraic structures give a generalisation of the structure of group, one removing the axiom about the existence of inverses and the other by relaxing the domain of the product (i.e. they can be tuples for which the product is not define). In these two categories, it is much more easier to find an object with a given number of automorphisms. More precisely, if $\mathcal{C}$ is the category ${{\mathcal{M}}onoids}$ or ${{\mathcal{P}}art}$, for every group $G$, one can find infinitely many objects of $\mathcal{C}$, with $G$ as automorphism group. This realisability result for partial group is due to the author in collaboration with Diaz and Viruel [6]. The corresponding one on monoids is not new and can be deduced from the well know construction highlighted in section 3.1 bellow, and theorems from Frucht [5] and Sabidussi [7] on the realisability for graphs which state that every group is the automorphism group of a graph. This is actually the same tactic that is use in [6]: we first prove that for every graph $\Gamma$ we can find a partial group $\mathcal{M}$ with the same automorphism group as $\Gamma$ and we conclude using the results of Frucht and Sabidussi.
Here, after giving functors from the category of graphs to ${{\mathcal{M}}onoids}$ and $Part$ which induce isomorphisms on automorphism groups, we consider well chosen graphs to obtain the followings.
Theorem 1.2 (cf Corollary 4.4).
Let $r\in\mathbb{Q}\cap]0,+\infty[$.
There exists a monoid $M$ such that $r=\dfrac{|\text{Aut}_{{\mathcal{M}}onoids}(M)|}{|M|}$. Moreover, the monoid $M$ can be chosen to be commutative.
Theorem 1.3 (cf Corollary 4.5).
Let $r\in\mathbb{Q}\cap]0,+\infty[$.
There exists a partial group $\mathcal{M}$ such that $r=\dfrac{|\text{Aut}_{{\mathcal{P}}art}(\mathcal{M})|}{|\mathcal{M}|}$.
Note that the construction used for partial groups is different from the one in [6]. Indeed, in [6], the partial groups considered are infinite even when the graph is finite. Hence the construction given in Section 3.2 leads to another proof of the universality of the category of partial groups.
We also use the same tactic on posets to get the same kind of result in that context.
Theorem 1.4 (cf Corollary 4.6).
Let $r\in\mathbb{Q}\cap]0,+\infty[$.
There exists a poset $P$ such that $r=\dfrac{|\text{Aut}_{{\mathcal{P}}osets}(P)|}{|P|}$.
Along the way, we also get a similar results for graphs where the ”size” of a graph is either the number of vertices or the number of vertices plus the number of edges.
Theorem 1.5 (cf Corollary 4.2 and Corollary 4.3).
Let $r\in\mathbb{Q}\cap]0,+\infty[$.
  1. (1)
    There exists a graphs $\Gamma$ such that $r=\dfrac{|\text{Aut}_{{\mathcal{G}}raphs}(\Gamma)|}{|V|}$ where $V$ is the set of vertices of $\Gamma$.
  2. (2)
    There exists a graphs $\Gamma$ such that $r=\dfrac{|\text{Aut}_{{\mathcal{G}}raphs}(\Gamma)|}{|V|+|E]}$ where $V$ and $E$ are the sets of, respectively, vertices and edges of $\Gamma$.
Notation
In these notes, if $\mathcal{C}$ is a category, for every object $x$ of $\mathcal{C}$, we will denote by $\text{Aut}_{\mathcal{C}}(x)$ the group of automorphisms of $x$ in $\mathcal{C}$. Also if $X$ is a finite set $|X|$ will denote the size of $X$. For $X$ a set $\mathbb{W}(X)$ will denote the free monoid of the finite words on $X$, where the elements will be denoted by a list of elements of $X$, and for two words $u,v\in\mathbb{W}(X)$, $u\circ v$ will denote the concatenation of $u$ and $v$. We often identify $X$ as the subset $\{(x)\mid x\in X\}$ of words of length 1 in $\mathbb{W}(X)$. Finally, given two sets $X$ and $Y$ and a map $\varphi\colon X\to Y$ we will denote by $\mathbb{W}(\varphi)\colon\mathbb{W}(X)\to\mathbb{W}(Y)$ the induce map by $\varphi$ defined by, for $u=(x_{1},x_{2},\dots,x_{n})\in\mathbb{W}(X)$, $\mathbb{W}(\varphi)(u)=(\varphi(x_{1}),\varphi(x_{2}),\dots,\varphi(x_{n}))$.
Finally, for $p\geq 1$, $\Sigma_{p}$ will denote the symmetric group on $\{1,2,\dots,p\}$ which is of order $p!$.
Acknowledgement
The author is thankful to Antonio Viruel for pointing out a mistake in the first draft of this paper.
2. recalls and notations on the algebraic structures
We recall here the algebraic structures and give the associated notations that we use in this paper.
2.1. Graphs
We work here with the category ${{\mathcal{G}}raphs}$ of graphs which are undirected and simple (without loops or multiple edges). More precisely, a graph is a couple $(V,E)$ with $V$ a set, called the set of vertices and $E$ a set of subsets of $V$ of size 2, called the set of edges. A vertex $x\in V$ will be called isolated if there is no edge of $E$ containing $x$. A graph homomorphism from a graph $\Gamma_{1}=(V_{1},E_{1})$ to a graph $\Gamma_{2}=(V_{2},E_{2})$ is then a map $f\colon V_{1}\to V_{2}$ such that, for any $e\in E_{1}$, $\overline{f}(e)\in E_{2}$ where $\overline{f}\colon\mathcal{P}(V_{1})\to\mathcal{P}(V_{2})$ is the mapping, induced by $f$, from $\mathcal{P}(V_{1})$ the set of all subsets of $V$ to $\mathcal{P}(V_{2})$ the set of all subsets of $V_{2}$. Finally, we will denote by ${\mathcal{I}nj{\mathcal{G}}raphs}$ the subcategory of ${{\mathcal{G}}raphs}$ with the same objects but with only injective graph homomorphisms. Notice that for any graph $\Gamma$, $\text{Aut}_{\mathcal{I}nj{\mathcal{G}}raphs}(\Gamma)=\text{Aut}_{{\mathcal{G}}raphs}(\Gamma)$ since every automorphism is injective.
2.2. Monoids
Recall that a monoid is a couple $(M,\ast)$ where $M$ is a set and $\ast\colon M\times M\to M$ is a binary operation which is associative and has a neutral element. A monoid homomorphism $\varphi\colon(M_{1},\ast_{1})\to(M_{2},\ast_{2})$ is a mapping such that for all $m,n\in M_{1}$, $\varphi(m\ast_{1}n)=\varphi(m)\ast_{2}\varphi(n)$ and such that $\varphi(e_{1})=e_{2}$ where $e_{1}$ and $e_{2}$ are the neutral elements of $M_{1}$ and $M_{2}$ respectively. With this notion of homomorphisms the class of monoids defines a category ${{\mathcal{M}}onoids}$. The category ${{\mathcal{G}}roups}$ of groups is the full subcategory of ${{\mathcal{M}}onoids}$ consisting of the monoids $(M,\ast)$ such that every element of $M$ has an inverse.
2.3. Partial groups
The notion of partial group is less common and we give slightly more details. We refer the reader to [3, Section 2] or [4, Section 1] for more comprehensive introductions about this structure.
Definition 2.1.
Let $\mathcal{M}$ be a set and $\mathbb{D}\subseteq\mathbb{W}(\mathcal{M})$ be a subset such that,
  1. (D1)
    $\mathcal{M}\subseteq\mathbb{D}$; and
  2. (D2)
    $u\circ v\in\mathbb{D}\Rightarrow u,v\in\mathbb{D}(\mathcal{M})$ (in particular, $\emptyset\in\mathbb{D}(\mathcal{M})$).
A mapping $\Pi:\mathbb{D}\rightarrow\mathcal{M}$ is a product if
  1. (P1)
    $\Pi$ restricts to the identity on $\mathcal{M}$; and
  2. (P2)
    if $u\circ v\circ w\in\mathbb{D}(\mathcal{M})$ then $u\circ\Pi(v)\circ w\in\mathbb{D}(\mathcal{M})$ and
    $\Pi(u\circ v\circ w)=\Pi\left(u\circ\Pi(v)\circ w\right).$
The unit of $\Pi$ is then defined as $\Pi(\emptyset)$ that we will denote $1_{\mathcal{M}}$ or $1$ when there is no ambiguities. A partial monoid is a triple $(\mathcal{M},\mathbb{D},\Pi)$ where $\Pi$ is a product defined on $\mathbb{D}(\mathcal{M})$.
An inversion on $\mathcal{M}$ is an involutory bijection $x\mapsto x^{-1}$ on $\mathcal{M}$ together with the induced mapping $u\mapsto u^{-1}$ on $\mathbb{W}(\mathcal{M})$ defined by,
$u=(x_{1},x_{2},\dots,x_{n})\mapsto(x_{n}^{-1},x_{n-1}^{-1},\dots,x_{1}^{-1}).$
A partial group is a tuple $\left(\mathcal{M},\mathbb{D},\Pi,(-)^{-1}\right)$ where $\Pi$ is a product on $\mathbb{D}(\mathcal{M})$ and $(-)^{-1}$ is an inversion on $\mathcal{M}$ satisfying
  1. (P3)
    if $u\in\mathbb{D}$ then $(u^{-1},u)\in\mathbb{D}$ and $\Pi(u^{-1}\circ u)=1$.
To simplify the notation, we will use $(\mathcal{M},\mathbb{D})$ or even juste $\mathcal{M}$ to refer to a partial group $\left(\mathcal{M},\mathbb{D},\Pi,(-)^{-1}\right)$ when the rest of the data is understood or not used. Also, a word $w\in\mathbb{D}$ will be called non-degenerate if there is no $1$ in $w$. Notice that $\mathbb{D}$ is totally determined by its non-degenerated word.
Example.
Let $(\mathcal{M},\mathbb{D})$ be a partial group. if the domain $\mathbb{D}=\mathbb{W}(\mathcal{M})$ then $\mathcal{M}$ is a group via the binary operation $(x,y)\in\mathcal{M}^{2}\mapsto\Pi(x,y)\in\mathcal{M}$.
Definition 2.2.
Let $\left(\mathcal{M}_{1},\mathbb{D}(\mathcal{M}_{1}),\Pi_{1},(-)^{-1}\right)$ and $\left(\mathcal{M}_{2},\mathbb{D}(\mathcal{M}_{2}),\Pi_{2},(-)^{-1}\right)$ be two partial group. A partial groups homomorphism is a map $\varphi\colon\mathcal{M}_{1}\to\mathcal{M}_{2}$ such that
  1. (H1)
    $\mathbb{W}(\varphi)\left(\mathbb{D}(\mathcal{M}_{1})\right)\subseteq\mathbb{D}(\mathcal{M}_{2})$;
  2. (H2)
    for any $u\in\mathbb{D}(\mathcal{M}_{1})$, $\Pi_{2}\left(\mathbb{W}(\varphi)(u)\right)=\varphi\left(\Pi_{1}(u)\right)$.
With this notion of homomorphisms and the usual composition on maps, we have a category ${{\mathcal{P}}art}$ which contains ${{\mathcal{G}}roups}$ as a full subcategory.
2.4. Posets
We will also work here with posets. Recall that a poset is a couple $(X,\leq)$ where $X$ is a set and $\leq$ a partial order, i.e. a homogeneous relation which is reflexive, antisymmetric, and transitive. A poset homomorphism $\varphi\colon(X_{1},\leq_{1})\to(X_{2}\leq_{2})$ is then a mapping $\varphi\colon X_{1}\to X_{2}$ which is order-preserving, i.e. for all $x,y\in X_{1}$, if $x\leq_{1}y$, then $\varphi(x)\leq_{2}\varphi(y)$. Endowed with this notion of morphism, the class of posets defines a category that we will denote by ${{\mathcal{P}}osets}$.
3. From graphs to Monoids, partial groups and posets
We give here the functors that we will use to prove the main results.
3.1. From graphs to monoids
The following construction is well known but the author couldn’t find a reference in the literature. However there is a discussion on math stack exchange [1] highlighting it.
Let $\Gamma=(V,E)$ be a graph. Set $M(\Gamma)$ be the union of $V$, $E$ and a arbitrary element denoted 00
$M(\Gamma)=V\cup E\cup\{0,1\}.$
We then define a product $\ast\colon M(\Gamma)\times M(\Gamma)\to M(\Gamma)$ setting, 00 as a absorbing element on $M(\Gamma)\setminus\{1\}$ (i.e. for all $\alpha\in M(\Gamma)\setminus\{1\}$, $00\ast\alpha=\alpha\ast 0=$ 0), $1$ as a neutral element (i.e. for all $\alpha\in M(\Gamma)$, $1\ast\alpha=\alpha\ast 1=\alpha$) and for all $x,x_{1},x_{2}\in V$, $e,e_{1},e_{2}\in E$
$x_{1}\ast x_{2}=\begin{cases}x_{1}&\text{if $x_{1}=x_{2}$}\\ 0&\text{else}\end{cases}\qquad\qquad x\ast e=e\ast x=\begin{cases}e&\text{if $x\in e$}\\ 0&\text{else,}\end{cases}\qquad\qquad e_{1}\ast e_{2}=0.$
Proposition 3.1.
Let $\Gamma=(V,E)$ be a graph. The couple $(M(\Gamma),\ast)$ is a monoid and it yields a functor
$M\colon{\mathcal{I}nj{\mathcal{G}}raphs}\to{{\mathcal{M}}onoids}.$
Proof.
The fact that $(M(\Gamma),\ast)$ is a monoid is a direct consequence of the definition. If $f\colon\Gamma_{1}=(V_{1},E_{1})\to\Gamma_{2}=(V_{2},E_{2})$ is a, injective graph homomorphism, it sends a vertex to a vertex, an edge to an edge and, for all $x\in V_{1}$ and $e\in E_{1}$, $x\in e_{1}$ if and only if $f(x)\in\overline{f}(e_{1})$. Thus it induces a well define monoid homomorphism. This construction preserves composition and thus defines a functor $M\colon{{\mathcal{G}}raphs}\to{{\mathcal{M}}onoids}$. ∎
The following Lemma is a direct consequence of the definition of $M(\Gamma)$.
Lemma 3.2.
Let $\Gamma=(V,E)$ be a graph.
  1. (i)
    $M(\Gamma)$ is commutative.
  2. (ii)
    For all $\alpha\in M(\Gamma)\setminus\{0,1\}$, $\alpha\ast\alpha=\alpha$ if and only if $\alpha\in V$.
  3. (iii)
    For all $e\in M(\Gamma)\setminus\{0\}$, $\alpha\ast\alpha=0$ if and only if $\alpha\in E$.
  4. (iv)
    For all $x\in V$ and $e\in E$, $x\ast e\neq 0$ if and only if $x\in e$.
  5. (v)
    If $\Gamma$ is finite then so is $M(\Gamma)$ and $|M(\Gamma)|=|V|+|E|+2$.
We now look at isomorphisms.
Lemma 3.3.
Let $\Gamma_{1}=(V_{1},E_{1})$ and $\Gamma_{2}=(V_{2},E_{2})$ be graphs and $\varphi\colon M(\Gamma_{1})\to M(\Gamma_{2})$ be an monoid isomorphism.
  1. (a)
    $\varphi(0)=0$, $\varphi(V_{1})=V_{2}$ and $\varphi(E_{1})=E_{2}$.
  2. (b)
    Let $x,y\in V_{1}\subseteq M(\Gamma_{1})$ with $x\neq y$. If $\{x,y\}\in E_{1}$ then $\{\varphi(x),\varphi(y)\}\in E_{2}$ and $\varphi(\{x,y\})=\{\varphi(x),\varphi(y)\}$.
Proof.
Notice that we have $\varphi(1)=1$ (as $\varphi$ is a monoid homomorphism) and $\varphi(0)=0$ since 00 is the only absorbing element on $M(\Gamma_{i})\setminus 1$ for $i\in\{1,2\}$. Moreover, a monoid homomorphism sends an idempotent to an idempotent, thus by Lemma 3.2(ii) we have $\varphi(V_{1})=V_{2}$ (considering $\varphi$ and its inverse). finally
$\varphi(E_{1})=\varphi\left(M(\Gamma_{1})\setminus\left(V_{1}\cup\{0,1\}\right)\right)=M(\Gamma_{2})\setminus\varphi\left(V_{1}\cup\{0,1\}\right)=M(\Gamma_{2})\setminus\left(V_{1}\cup\{0,1\}\right)=E_{2}$
which conclude the proof of item (a).
Now let $x,y\in V_{1}\subseteq M(\Gamma_{1})$ with $x\neq y$ such that $e=(x,y)\in E_{1}$. by (a), $\varphi(e)\in E_{2}$. Moreover, since $x\ast e=e$ and $y\ast e=e$, we have that $\varphi(x)\ast\varphi(e)=\varphi(e)$ and $\varphi(y)\ast\varphi(e)=\varphi(e)$, and by Lemma 3.2(iv), $\{\varphi(x),\varphi(y)\}\subseteq\varphi(e)$. Finally, since $\varphi$ is injective and $x\neq y$, we have $\varphi(x)\neq\varphi(y)$ and thus $\varphi(e)=\{\varphi(x),\varphi(y)\}$. ∎
Theorem 3.4.
Let $\Gamma=(V,E)$ be a graph. Then $\text{Aut}_{{\mathcal{G}}raphs}(\Gamma)\cong\text{Aut}_{{\mathcal{M}}onoids}(M(\Gamma))$.
Proof.
Since $M\colon{\mathcal{I}nj{\mathcal{G}}raphs}\to{{\mathcal{M}}onoids}$ is a functor, $M$ induces a group homomorphism $\Phi\colon\text{Aut}_{{\mathcal{G}}raphs}(\Gamma)=\text{Aut}_{\mathcal{I}nj{\mathcal{G}}raphs}(\Gamma)\to\text{Aut}_{{\mathcal{M}}onoids}(M(\Gamma))$. Moreover, notice that $f=M(f)|_{V}$ and thus, if $M(f)$ is the identity, so is $f$. Hence $\Phi$ is injective. Finally, by Lemma 3.3, if $\varphi\in\text{Aut}_{{\mathcal{M}}onoids}(M(\Gamma))$, then $\varphi|_{V}$ defines a graph homomorphism from $\Gamma$ to itself which is an automorphism of $\Gamma$ and we have $\Phi(\varphi|_{V})=\varphi$. Hence $\Phi$ is an isomorphism. ∎
3.2. From graphs to partial groups
Let $\Gamma=(V,E)$ be a graph. Set $\mathcal{E}(\Gamma)$ as the subset of $\mathbb{W}(V)$ consisting in the empty word together with the words of length 1 and the words of length 2 corresponding to edges in $E$ :
$\mathcal{E}(\Gamma)=\{\emptyset\}\cup\{(x)\mid x\in V\}\cup\{(x,y)\mid\{x,y\}\in E\}\subset\mathbb{W}(V).$
The domain $\mathbb{D}(\Gamma)$ will be define as the set of words $u=(u_{1},u_{2},\dots,u_{r})$ so that the concatenation $u_{1}\circ u_{2},\circ\cdots\circ u_{r}$ is a word of the form $(x,x,\dots,x)$ for $x\in V$ (in particular, for all $i\in\{1,2\dots,r\}$, $u_{i}=(x)$ or $u_{i}=\emptyset$) or a finite subword of an infinite word of the form
$(\underbrace{x,x,\dots,x}_{2k_{1}\text{ terms}},\underbrace{y,y,\dots,y}_{2r_{1}\text{ terms}},\underbrace{x,x,\dots,x}_{2k_{2}\text{ terms}},\underbrace{y,y,\dots,y}_{2r_{2}\text{ terms}},\dots)$
with $\{x,y\}\in E$ and for all $j$, $k_{j},r_{j}\geq 1$. For example, if $\{x,y\}\in E$, $(x,x,x,y,y,y,y,y,x,x,y,y,y,y,y,y,x)\in\mathbb{D}(\Gamma)$.
The product $\Pi\colon\mathbb{D}(\Gamma)\to\mathcal{E}(\Gamma)$ is then given, for $u=(u_{1},u_{2},\dots,u_{r})\in\mathbb{D}(\Gamma)$, by $\Pi(u)$ as the word obtained from the concatenation $u_{1}\circ u_{2},\circ\cdots\circ u_{r}$ by removing subwords of the form $(x,x)$ for $x\in V$ until no such reduction is possible (one can check that the resulting word is independent of the reduction process). In particular, since by definition of $\mathbb{D}(\Gamma)$ $u$ involves at most two vertices, and two vertices of the same edge when it involves exactly two, we have that $\Pi(u)\in\mathcal{E}(\Gamma)$. Notice also that the empty word plays here the role of the unit.
Finally, the inversion $(-)^{-1}$ is defined by sending a word $(x)$ to its symmetric words (i.e. for all $x\in V$, $(x)^{-1}=(x)$ and for all $\{x,y\}\in E$, $(x,y)^{-1}=(y,x)$).
Proposition 3.5.
Let $\Gamma=(V,E)$ be a graph. the tuple $\left(\mathcal{E}(\Gamma),\mathbb{D}(\Gamma),\Pi,(-)^{-1}\right)$ constructed above defines a partial group. Moreover, this construction gives a functor $\mathcal{E}\colon{{\mathcal{G}}raphs}\to{{\mathcal{P}}art}$.
Proof.
By construction of $\mathbb{D}(\Gamma)$, the conditions (D1) and (D2) of Definition 2.1 are satisfied. By construction of $\Pi$, the condition (P1) is satisfy. For condition (P2) a case-by-case study shows that it is also satisfied. For example, let $u,v,w\in\mathbb{D}(\Gamma)$ and assume that $u\circ v\circ w\in\mathbb{D}(\Gamma)$ is a word, involving two vertices $x,y\in V$. Assume also that $u,v$ and $w$ are non empty and that $v$ is of the form
$(x,\underbrace{x,x,\dots,x}_{2k_{1}\text{ terms}},\underbrace{y,y,\dots,y}_{2r_{1}\text{ terms}},\underbrace{x,x,\dots,x}_{2k_{2}\text{ terms}},\underbrace{y,y,\dots,y}_{2r_{2}\text{ terms}},\dots,\underbrace{x,x,\dots,x}_{2k_{n}\text{ terms}},\underbrace{y,y,\dots,y}_{2r_{n}\text{ terms}})$
with $n\geq 1$ and for all $j\in\{1,2,\dots,n\}$, $k_{j},r_{j}\geq 1$. Then, since $u$ is not empty, by definition of $\mathbb{D}(\Gamma)$ we can write $u=u^{\prime}\circ(x)$ with $u^{\prime}\in\mathbb{D}(\Gamma)$. By definition of $\Pi$, we get that
$\Pi(u\circ v\circ w)$ $=\Pi(u^{\prime}\circ(x)\circ v\circ w)$
$=\Pi(u^{\prime}\circ(\underbrace{x,x,\dots,x}_{2k_{1}+2\text{ terms}},\underbrace{y,y,\dots,y}_{2r_{1}\text{ terms}},\dots,\underbrace{x,x,\dots,x}_{2k_{n}\text{ terms}},\underbrace{y,y,\dots,y}_{2r_{n}\text{ terms}})\circ w)$
$=\Pi(u^{\prime}\circ w).$
We also have $\Pi(v)=(x)$ and thus $\Pi(u\circ\Pi(v)\circ w)=\Pi(u^{\prime}\circ(x)\circ(x)\circ w)=\Pi(u\circ(x,x)\circ w)=\Pi(u^{\prime}\circ w)$. In particular, $\Pi(u\circ v\circ w)=\Pi(u\circ\Pi(v)\circ w)$. The other cases can be managed in a similar fashion. Finally, the definition of the inversion, $\mathbb{D}(\Gamma)$ and $\Pi$ ensure that condition (P3) is satisfied.
For the functoriality, notice that, if $\Gamma_{1}=(V_{1},E_{1})$ and $\Gamma_{2}=(V_{2},E_{2})$ are two graphs and $f\colon\Gamma_{1}\to\Gamma_{2}$ is a graph homomorphism, $f$ induces a natural map $\mathbb{W}(f)\colon\mathbb{W}(V_{1})\to\mathbb{W}(V_{2})$ which sends $\mathcal{E}(\Gamma_{1})$ to $\mathcal{E}(\Gamma_{2})$. This restriction gives a partial groups homomorphism $\mathcal{E}(f)\colon\mathcal{E}(\Gamma_{1})\to\mathcal{E}(\Gamma_{2})$ and all this is compatible with composition of graph homomorphisms. ∎
The following Lemma is a direct consequence of the partial group structure on $\mathcal{E}(\Gamma)$ given above.
Lemma 3.6.
Let $\Gamma=(V,E)$ be a graph.
  1. (i)
    For all $u\in\mathcal{E}(\Gamma)$, $(u,u)\in\mathbb{D}(\Gamma)$ if and only if there exists $x\in V$ such that $u=(x)$.
  2. (ii)
    For all $x,y\in V$ with $x\neq y$, $((x),(y))\in\mathbb{D}(\Gamma)$ if and only if $\{x,y\}\in E$. Moreover $(x,y)=\Pi((x),(y))$.
  3. (iii)
    If $\Gamma$ is finite, then $\mathcal{E}(\Gamma)$ is also finite and $|\mathcal{E}(\Gamma)|=1+|V|+2|E|$.
We now look more closely at what happens on morphisms.
Lemma 3.7.
Let $\Gamma_{1}=(V_{1},E_{1})$ and $\Gamma_{2}=(V_{2},E_{2})$ be two finite graphs and $\varphi\colon\mathcal{M}(\Gamma_{1})\to\mathcal{M}(\Gamma_{2})$ a partial group homomorphism.
  1. (a)
    $\varphi$ is totally determined by its image on $\{(x)\mid x\in V_{1}\}\subseteq\mathcal{E}(\Gamma_{1})$.
  2. (b)
    $\varphi\left(\{(x)\mid x\in V_{1}\}\right)\subseteq\{(x)\mid x\in V_{2}\}$.
  3. (c)
    Assume $\varphi$ is injective and let $x,y\in V_{1}$ with $x\neq y$. If $\{x,y\}\in E_{1}$, then $\{\overline{\varphi}(x),\overline{\varphi}(y)\}\in E_{2}$ and $\varphi((x,y))=(\overline{\varphi}(x),\overline{\varphi}(y))$, where $\overline{\varphi}\colon V_{1}\to V_{2}$ is the application such that, for all $x\in V_{1}$, $\varphi((x))=((\overline{\varphi}(x))$ ($\overline{\varphi}$ exists by (b)).
Proof.
Part (a) follows from Lemma 3.6(ii) since $\varphi$ is a partial group homomorphism.
By Lemma 3.6(i), for all $i\in\{1,2\}$, the subset $\{(x)\mid v\in V_{i}\}\subseteq\mathcal{E}(\Gamma_{i})$ is exactly the set of all the non trivial elements of $\mathcal{E}(\Gamma_{i})$ that can be multiplied with themselves. Therefore $\varphi(\{(x)\mid x\in V_{1}\})\subseteq\{(x)\mid x\in V_{2}\}$ and (b) follows.
Finally, assume that $\varphi$ is injective and let $x,y\in V_{1}$. Assume moreover that $\{x,y\}\in E_{1}$. Then, $(x,y)\in\mathcal{E}(\Gamma_{1})$ and by Lemma 3.6(ii) we have $((x),(y))\in\mathbb{D}(\Gamma_{1})$ and $(x,y)=\Pi((x),(y))$. Therefore $((\overline{\varphi}(x)),(\overline{\varphi}(y)))=(\varphi(x),\varphi(y))\in\mathbb{D}(\Gamma_{2})$. Now since $\varphi$ is injective, $\overline{\varphi}(x)\neq\overline{\varphi}(y)$ and, again by Lemma 3.6(ii), $\{\overline{\varphi}(x),\overline{\varphi}(y)\}\in\mathcal{E}(\Gamma_{2})$ and $\Pi((\overline{\varphi}(x)),(\overline{\varphi}(y)))=(\overline{\varphi}(x),\overline{\varphi}(y))$. Finally, since $\varphi$ is a partial group homomorphism, $\varphi((x,y))=\varphi(\Pi((x),(y)))=\Pi(\varphi((x)),\varphi((y)))=\Pi((\overline{\varphi}(x)),(\overline{\varphi}(y)))=(\overline{\varphi}(x),\overline{\varphi}(y))$. ∎
Theorem 3.8.
Let $\Gamma=(V,E)$ be a graph. Then $\text{Aut}_{{\mathcal{G}}raphs}(\Gamma)=\text{Aut}_{{\mathcal{P}}art}\left(\mathcal{E}(\Gamma)\right)$.
Proof.
By functoriality, $\mathcal{E}\colon{{\mathcal{G}}raphs}\to{{\mathcal{P}}art}$ induces a group morphism $\Phi\colon\text{Aut}_{{\mathcal{G}}raphs}(\Gamma)\to\text{Aut}_{{\mathcal{P}}art}\left(\mathcal{E}(\Gamma)\right)$.
Now let $f$ be a automorphism of $\Gamma$ and assume that $\Phi(f)=\text{Id}$. In particular, by Lemma 3.7(b), for every $x\in V$, $f(x)=\overline{\Phi(f)}(x)=x$. Thus $f=\text{Id}_{\Gamma}$.
Finally, let $\varphi\in\text{Aut}_{{\mathcal{P}}art}(\mathcal{E}(\Gamma))$ and set $\mathcal{E}_{V}=\{(x)\mid x\in V\}$. By Lemma 3.7(b) and since $\varphi$ is an automorphism, for all $x\in V$, $\varphi((x))\in\mathcal{E}_{V}$ and f induces a bijection $\overline{\varphi}\colon V\to V$. Then, by Lemma 3.7(c), $\overline{\varphi}\in\text{Aut}_{{\mathcal{G}}raphs}(\Gamma)$. Moreover, $\Phi(\overline{\varphi})=\varphi$ and so $\Phi$ is surjective. ∎
3.3. From graph to posets
Finally, we construct a functor from ${{\mathcal{G}}raphs}$ to ${{\mathcal{P}}osets}$. What follows is not really knew and was already highlighted in the introduction of [2].
Let $\Gamma=(V,E)$ be a graph, we consider the poset $P(\Gamma)=\{\{x\}\mid x\in V\}\cup E$, i.e. $P(G)$ is a subset of the set of subsets of $V$ consisting in all the singleton in $V$ and the elements of $E$ (which are subset of $V$ of size 2), order by inclusion. This poset is sometimes called the face poset of $\Gamma$ and it defines a functor $P\colon{{\mathcal{G}}raphs}\to{{\mathcal{P}}osets}$.
The following Lemma is a direct consequence of the construction.
Lemma 3.9.
Let $\Gamma=(V,E)$ be a graph. We will denote by $\leq$ the partial order on $P(\Gamma)$
  1. (i)
    The set of maximal elements of $P(\Gamma)$ is $E$ together with the isolated vertices of $\Gamma$ and the set of minimal element of $P(\Gamma)$ is $V$.
  2. (ii)
    For all $u,v\in P(\Gamma)$ with $u\neq v$, $u\leq v$ if and only if $v\in E$ and there exists $x\in V$ such that $u=\{x\}$ and $x$ is one of the extremities of $v$.
  3. (iii)
    If $\Gamma$ is a finite, then $P(\Gamma)$ is also finite and $|P(\Gamma)|=|V|+|E|$.
Let us now look at what happens on isomorphisms.
Lemma 3.10.
Let $\Gamma_{1}=(V_{1},E_{1})$ and $\Gamma_{2}=(V_{2},E_{2})$ be two finite graphs and $\varphi\colon P(\Gamma_{1})\to P(\Gamma_{2})$ a poset homomorphism and assume that $\varphi$ is an isomorphism.
  1. (a)
    $\varphi(\{\{x\}\mid x\in V_{1}\})=\{\{x\}\mid x\in V_{2}\}$ and $\varphi(E_{1})=E_{2}$.
  2. (b)
    Let $x,y\in V_{1}$ with $x\neq y$. If $\{x,y\}\in E_{1}$, then $\{\overline{\varphi}(x),\overline{\varphi}(y)\}\in E_{2}$ and $\varphi(\{x,y\})=\{\overline{\varphi}(x),\overline{\varphi}(y)\}$, where $\overline{\varphi}\colon V_{1}\to V_{2}$ is the application such that, for all $x\in V_{1}$, $\varphi(\{x\})=\{\overline{\varphi}(x)\}$.
Proof.
Since $\varphi$ is an isomorphism, then it sends the minimal, resp. maximal, elements of $P(\Gamma_{1})$ to the the minimal, resp. maximal, elements of $P(\Gamma_{2})$. Therefore, part (a) is consequence of Lemma 3.9(i).
Now, let $x,y\in V_{1}$ with $x\neq y$. If $\{x,y\}\in E_{1}$, by (a), $\varphi(\{x,y\})\in E_{2}$. So there exists $a,b\in V_{2}$ such that $\{a,b\}\in E_{2}$ and $\varphi(\{x,y\}=\{a,b\}$. Since $\{x\}\leq\{x,y\}$ and $\{y\}\leq\{x,y\}$ and $\varphi$ is a poset homomorphism, $\varphi(\{x\})=\{\overline{\varphi}(x)\}\subset\{a,b\}$ and $\varphi(\{y\})=\{\overline{\varphi}(y)\}\subset\{a,b\}$. in particular, by 3.9(ii), $\overline{\varphi}(x)\in\{a,b\}$ and $\overline{\varphi}(y)\in\{a,b\}$. As $\varphi$ is injective and $x\neq y$, we have $\{\overline{\varphi}(x),\overline{\varphi}(y)\}=\{a,b\}$ and part (b) follows. ∎
Theorem 3.11.
Let $\Gamma=(V,E)$ be a graph. Then $\text{Aut}_{{\mathcal{G}}raphs}(\Gamma)=\text{Aut}_{{\mathcal{P}}osets}\left(P(\Gamma)\right)$.
Proof.
By functoriality, the functor $P\colon{{\mathcal{G}}raphs}\to{{\mathcal{P}}osets}$ induces a group morphism $\Phi\colon\text{Aut}_{{\mathcal{G}}raphs}(\Gamma)\to\text{Aut}_{{\mathcal{P}}osets}\left(P(\Gamma)\right)$.
Now, let $f$ be a automorphism of $\Gamma$ and assume that $\Phi(f)=\text{Id}$. By Lemma 3.10(a), $f=\overline{\Phi(f)}=\text{Id}$ and so $\Phi$ is injective.
Finally, let $\varphi\in\text{Aut}_{{\mathcal{P}}osets}(\mathcal{E}(\Gamma))$ and set $P_{V}=\{(x)\mid x\in V\}$. By Lemma 3.10(a) f induces a bijection $\overline{\varphi}\colon V\to V$ and, by Lemma 3.10(b), $\overline{\varphi}\in\text{Aut}_{{\mathcal{G}}raphs}(\Gamma)$. Moreover, $\Phi(\overline{\varphi})=\varphi$ and so $\Phi$ is surjective. ∎
4. Some particular graphs
Fix $p,q\geq 1$ and set $\Gamma_{p,q}=(V_{p,q},E_{p,q})$ where $V_{p,q}$ is a set with $p+q+1$ elements denoted
$V_{p,q}=\{c_{0},c_{1},c_{2},\dots,c_{p},d_{1},d_{2},\dots,d_{q}\},$
and where
$E_{p,q}=\{\{(c_{0},c_{i}\}\mid i\in\{1,2,\dots,p\}\}\cup\{\{c_{0},d_{1}\}\}\cup\{\{d_{i},d_{i+1}\}\mid i\in\{1,2,\dots,q-1\}\}.$
For example here are $\Gamma_{2,3}$ and $\Gamma_{3,2}$.
$c_{1}$$c_{2}$$c_{0}$$d_{1}$$d_{2}$$d_{3}$$c_{1}$$c_{2}$$c_{3}$$c_{0}$$d_{1}$$d_{2}$
Finally we set $\Gamma_{p,q}^{+}=\Gamma_{p,q}\cup\{c_{0}^{\prime}\}$ to be the graph $\Gamma_{p,q}$ together with a single vertex added (and no edges).
Proposition 4.1.
Let $p,q\geq 1$, then $|V_{p,q}|=p+q+1$ and $|E_{p,q}|=p+q$. If $p,q\geq 2$ then the group $\text{Aut}_{{\mathcal{G}}raphs}(\Gamma_{p,q})\cong\text{Aut}_{{\mathcal{G}}raphs}(\Gamma_{p,q}^{+})\cong\Sigma_{p}$.
Proof.
The cardinalities of the vertex sets are a direct consequence of the definition. Notice also, since $c_{0}^{\prime}$ is the unique isolated vertex in $\Gamma_{p,q}^{+}$, it is fixed by any automorphism and so $\text{Aut}_{{\mathcal{G}}raphs}(\Gamma_{p,q}^{+})\cong\text{Aut}_{{\mathcal{G}}raphs}(\Gamma_{p,q})$. It remains to prove that $\text{Aut}_{{\mathcal{G}}raphs}(\Gamma_{p,q})\cong\Sigma_{p}$
Notice that $c_{0}$ is the only vertex of degree 3 or more. Hence it is fixed by any automorphism. Then the tail $c_{0}-d_{1}-d_{2}-\cdots-d_{q}$, which is the only tail starting at $c_{0}$ of length 2 or more, is also fixed by any automorphism. Therefore, $\text{Aut}_{{\mathcal{G}}raphs}(\Gamma_{p,q})$ acts faithfully on $\{c_{1},c_{2},\dots,c_{p}\}$ and this action gives an isomorphism $\text{Aut}_{{\mathcal{G}}raphs}(\Gamma_{p,q})\cong\Sigma_{p}$. ∎
In particular the ratio $\dfrac{|\text{Aut}_{{\mathcal{G}}raphs}(\Gamma_{p,q})|}{|V_{p,q}|}$ is equal to $\dfrac{p!}{p+q+1}$ and we have the following corollary.
Corollary 4.2.
Let $r\in\mathbb{Q}\cap]0,+\infty[$. There exists a graph $\Gamma=(V,E)$ such that $r=\dfrac{|\text{Aut}_{{\mathcal{G}}raphs}(\Gamma)|}{|V|}$.
Proof.
take $m,n\geq 3$ such that $r=\dfrac{m}{n}$. Then taking $p=m$ and $q=n\times(m-1)!-m-1$, we have $p\geq 2$, $q\geq 3(m-1)-m-1=2m-4\geq 2$ and thus
$\frac{|\text{Aut}_{{\mathcal{G}}raphs}(\Gamma_{p,q})|}{|V_{p,q}|}=\frac{p!}{p+q+1}=\frac{m!}{n\times(m-1)!}=\frac{m}{n}=r.$
Now if we consider the graph $\Gamma_{p,q}^{+}$ for $p$ and $q$ well chosen to get the following two corollaries.
Corollary 4.3.
Let $r\in\mathbb{Q}\cap]0,+\infty[$. There exists a graph $\Gamma=(V,E)$ such that $r=\dfrac{|\text{Aut}_{{\mathcal{G}}raphs}(\Gamma)|}{|V|+|E]}$.
Proof.
take $m\geq 3$ and $n\geq 2$ such that $r=\dfrac{m}{n}$. Then taking $p=2m$ and $q=n\times(2m-1)!-2m-1$, we have $p\geq 6>2$, $q\geq 2(2m-1)-2m-1=2m-3>2$ and thus the wanted quotient is,
$\frac{|\text{Aut}_{{\mathcal{G}}raphs}(\Gamma_{p,q}^{+})|}{|V_{p,q}|+1+|E_{p+q}|}=\frac{p!}{2(p+q+1)}=\frac{(2m)!}{2\times n\times(2m-1)!}=\frac{2m}{2n}=r.$
Corollary 4.4.
Let $r\in\mathbb{Q}\cap]0,+\infty[$. There exists a monoid $M$ such that $r=\dfrac{|\text{Aut}_{{\mathcal{M}}onoids}(M)|}{|M|}$. Moreover, the monoid $M$ can be chosen to be commutative.
Proof.
take $m\geq 3$ and $n\geq 2$ such that $r=\dfrac{m}{n}$. Then taking $p=2m$ and $q=n\times(2m-1)!-2m-2$, we have $p\geq 6>2$, $q\geq 2(2m-1)-2m-2=2m-4\geq 2$ and thus, by Theorem 3.4 and Proposition 4.1,
$|\text{Aut}_{{\mathcal{M}}onoids}(M(\Gamma_{p,q}))|=|\text{Aut}_{{\mathcal{G}}raphs}(\Gamma_{p,q})|=p!=(2m)!.$
Also by Lemma 3.2(v),
$|M(\Gamma_{p,q})|=|V_{p,q}|+|E_{p,q}|+2=2(p+q)+4=2n\times(2m-1)!$
Therefore,
$\frac{|\text{Aut}_{{\mathcal{M}}onoids}(M(\Gamma_{p,q}))|}{|M(\Gamma_{p,q})|}=\frac{(2m)!}{2n\times(2m-1)!}=\frac{2m}{2n}=\frac{m}{n}=r.$
Finally, recall that by Lemma 3.2(i), $M$ is commutative. ∎
Corollary 4.5.
Let $r\in\mathbb{Q}\cap]0,+\infty[$. There exists a partial group $\mathcal{M}$ such that $r=\dfrac{|\text{Aut}_{{\mathcal{P}}art}(\mathcal{M})|}{|\mathcal{M}|}$.
Proof.
take $m,n\geq 2$ such that $r=\dfrac{m}{n}$. Then taking $p=3m$ and $q=n\times(3m-1)!-3m-1$, we have $p\geq 6>2$ and $q\geq 2(3m-1)-3m-1=3m-3>2$ and thus , by Theorem 3.8 and Proposition 4.1,
$|\text{Aut}_{{\mathcal{P}}art}(\mathcal{E}(\Gamma_{p,q}^{+})|=|\text{Aut}_{{\mathcal{G}}raphs}(\Gamma_{p,q}^{+})|=p!=(3m)!.$
Also by Lemma 3.6(iii),
$|\mathcal{E}(\Gamma_{p,q}^{+}))|=(|V_{p,q}|+1)+2|E_{p,q}|+1=(p+q+2)+2(p+q)+1=3(p+q+1)=3n\times(3m-1)!.$
Therefore,
$\frac{|\text{Aut}_{{\mathcal{P}}art}(\mathcal{E}(\Gamma_{p,q}^{+}))|}{|\mathcal{E}(\Gamma_{p,q}^{+})|}=\frac{(3m)!}{3\times n\times(3m-1)!}=\frac{m}{n}=r.$
Corollary 4.6.
Let $r\in\mathbb{Q}\cap]0,+\infty[$. There exists a poset $P$ such that $r=\dfrac{|\text{Aut}_{{\mathcal{P}}osets}(P)|}{|P|}$.
Proof.
take $m\geq 3$ and $n\geq 2$ such that $r=\dfrac{m}{n}$. Then taking $p=2m$ and $q=n\times(2m-1)!-2m-1$, we have $p\geq 6>2$, $q\geq 2(2m-1)-2m-1=2m-3\geq 3>2$ and thus, using Theorem 3.11, Proposition 4.1, Lemma 3.9(iii),
$\dfrac{|\text{Aut}_{{\mathcal{P}}osets}(P(\Gamma_{p,q}^{+}))|}{|P(\Gamma_{p,q}^{+})|}=\frac{|\text{Aut}_{{\mathcal{G}}raphs}(\Gamma_{p,q}^{+})|}{|P(\Gamma_{p,q}^{+})|}=\frac{p!}{2(p+q)+2}=\frac{2m!}{2n\times(2m-1)!}=\frac{2m}{2n}=\frac{m}{n}=r.$